Physics Problem

Suppose a straight 1.50 mm -diameter copper wire could just "float" horizontally in air because of the force due to the Earth's magnetic field B, which is horizontal, perpendicular to the wire, and of magnitude 4.1×10−5 T . Find the current the wire would carry.
Answer #1

a) let the current needed is I

Now,density of copper, p = 8960 Kg/m^3

Now, for the wire to levitate

lamda * g = B * I

8960 * pi * (0.0015/2)^2 = 5.5 *10^-5 * I

solving for I

I = 287 A

the current needed is 287 A

b) No

c) as this current would causes too high heat losses


answered by: ANURANJAN SARSAM